The area of a triangle is 40 sq. inches. If the base of the triangle is 10 inches, what is its height?
Answer by alicealc(293)

Area of triangle:
Area = 1/2 * base * height
40 = 1/2 * 10 * height
40 = 5 * height
height = 40/5 = 8

so, the triangle's height is 8 inches